OpenStudy (anonymous):

another function question Given the function f(x) = -2x + 8, find x if f(x) = 14

OpenStudy (anonymous):

-2(14) + 8 = -28 + 8 = ?

OpenStudy (anonymous):

its -20

OpenStudy (anonymous):

-20

OpenStudy (anonymous):

gj